OpenStudy (anonymous):

Let the function f(x) be defined by the equation y = f(x), where x and f(x) are real numbers. f(x) = 4 − x^2 f(9) = f(-7) = f(k) = f(k2 − 8) =

OpenStudy (anonymous):

f(9)=-77 f(-7)=-45 f(k)=4-k^2

OpenStudy (anonymous):

i'm not sure about the last one, sorry

OpenStudy (anonymous):

no worries.. how did you get -77? I want to try and understand this

OpenStudy (anonymous):

Just put in 9 to the equation so f(9)=4-9^2

OpenStudy (anonymous):

Last one is 2k-8? or k^2-8?

OpenStudy (anonymous):

i got it.. i was trying to multiply it instead of subtracting it...

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

andras 2k-8 nor k^2-8 are correct for the final answer ive tried them

OpenStudy (anonymous):

hmm, but my question was what is k2 here: f(k2 − 8)

OpenStudy (anonymous):

\[ f(k^2 − 8) = \] is the equation

OpenStudy (anonymous):

u just have to put k^2-8 into the function. So it will be f(k^2-8)= 4-(k^2-8)^2=-k^2+16k-60

OpenStudy (anonymous):

and I guess k is just a free parameter here